After doing my calculations, I determined Z=2.04. How were you able to determine that the test was one tailed? After making this determination, is this why the formula 1-p(z<2.04) was used. Is it because it was one tailed?
Answer #1

solution:

yes Is it because it was one tailed

this is to the right p(z>2.04) if this was goes to the left it will be 1-p(z<2.04)  

if z value is positive we must do this 1-p(z<2.04)  

p(z>2.04)=1-p(z<2.04) =1-0.9793=0.0207